Bentoods law first nou zero numher in a snolom distrihatin of its 56 PlXex)] 0.301 0.176 0.125 0.097 0,079 0.067 0.056 0.
Answer #1
observed frequencey, O expected proportion expected frequency,E (O-E)²/E
83 0.301 82.78 0.001
49 0.176 48.40 0.007
32 0.125 34.38 0.164
22 0.097 26.68 0.819
25 0.079 21.73 0.494
18 0.067 18.43 0.010
13 0.058 15.95 0.546
17 0.051 14.025 0.631
16 0.046 12.65 0.887

chi square test statistic,X² = Σ(O-E)²/E =   3.559             
                  
level of significance, α=   0.05              
Degree of freedom=k-1=   9   -   1   =   8
                  
P value =   0.8946   [ excel function: =chisq.dist.rt(test-stat,df) ]          
Decision: P value >α , Do not reject Ho                  

so, sample is according to law

......................


THANKS

revert back for doubt

please upvote